BRUHBRO:

1) x + y -19 and x - y = -3 solve by substitution

simsharrison:

4

Kiles14:

yeah

BRUHBRO:

how yk tho?

Kiles14:

work it out :)

BRUHBRO:

;-;

simsharrison:

im smart

BRUHBRO:

tbh idek how this problem works

Kiles14:

ok you on a lap top?

BRUHBRO:

yes

Kiles14:

ok go to chrome store, and download TI-84 plus CE app (its a calculator) dm me when you got it

BRUHBRO:

alrighty then

BRUHBRO:

breh

BRUHBRO:

i search it up and only some mario and dbz thing show up but no calculator

Kiles14:

smh look up calculator

BRUHBRO:

ok

BRUHBRO:

i dont see the ti-84 one

Kiles14:

oop

Kiles14:

mb i forgot smh

justus:

It has no solution.

BRUHBRO:

why not?

justus:

Actually, did you mean x+y=19 or x+y-19?

justus:

Sorry for the confusion.

BRUHBRO:

it dont matter no more

justus:

x+y=19 X-y=-3 19-y-y=-3 19-2y=-3 x=19-11 x=8 x=8, y=11

BRUHBRO:

i have the notes 😏

justus:

Okay good luck. :)

BRUHBRO:

wow you were still right

BRUHBRO:

*gives medal*

justus:

Thank you.

BRUHBRO:

lmao

Kiles14:

yeah never listen to me lmao